OpenStudy (anonymous):

(please help :( <3 I'll give a medal and a fan) 8. The Ordered pairs (1, 2), (2,4), (3, 8), (4, 16), and (5, 32) represent a function. What is a rule that represents this function? a. y= x^2 b.y=2x c. y= 2x y= x+2

OpenStudy (anonymous):

@triciaal @mathgeek27 helppp D:

OpenStudy (triciaal):

y = 2^x

OpenStudy (triciaal):

when x = 1 2^1 = 2 (1, 2) when x = 2 2^2 = 4 (2, 4) etc

OpenStudy (anonymous):

thank you :)